Session: May 1999 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Mon Jun 7 12:07:34 METDST 1999) Success
no problem happened

Westerbork (Tony Foley at Mon Jun 7 14:14:19 METDST 1999) Success
No known problems. Used 13 telescopes in tied array. Decoder channel B giving MUCH higher parity error rates than channel A.

Onsala (A. Polatidis at Mon Jun 7 16:50:47 METDST 1999) Minor failures
Telescope was stowed between 1317 and 1639 UT, due to high winds. Otherwise no problem with the data.

Jodrell (Alastair Gunn at Mon Jun 7 17:01:41 METDST 1999) Success
No known problems.

Cambridge (Alastair Gunn at Mon Jun 7 17:02:03 METDST 1999) Minor failures
Schedule not started on time due to operator error. Lost first 2 scans of data. No other known problems.

Effelsberg (A. Kraus at Tue Jun 8 08:30:39 METDST 1999) Minor failures
Telescope stop between 1410 - 1550 due to strong wind

Torun (K. Borkowski at Tue Jun 8 09:51:41 METDST 1999) Minor failures
Towards the end of the 3rd tape (after 4 UTC) the operator noticed that the second indicator of the formatter time is 1s ahead of UTC. This was corrected before the last tape start (4:50 UTC).

Noto (C.Trigilio at Tue Jun 8 12:06:38 METDST 1999) Minor failures
Hardware problem at the telescope from 08:34 UT. Telescope not working till end of experiment.

Calibration (Jean\-Francois Desmurs at Mon Jun 21 16:07:30 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the jun99 directory under the name: ec010.antab. Read the comments in the file to get further information. (Rem : Cm no Tsys values.)